Are the following lines parallel, perpendicular?
L1 with equation x – 2y = 4
L2 with equation 2x + y = 2

Answer by stanbon(75887)

Are the following lines parallel, perpendicular?
L1 with equation x – 2y = 4
slope-intercept form: y = (1/2)x -2 ; slope = 1/2
------------------------------------------------------------
L2 with equation 2x + y = 2
slope-intercept form: y = -2x + 2 ; slope = -2
------------------------------------------------------------
Since -2(1/2) = -1, the lines are perpendicular.
=========================================================
Cheers,
Stan H.